OpenStudy (anonymous):

atomic weight of two element (a) and (b) are 32 and 16 respectively they combine to give two gases ab2 and ab4 and the mixture is found to have a vapour density of 38.4 the number of moles of ab2 in 50 moles of the mixture is ??

OpenStudy (inkyvoyd):

Ok. The very first step is to realize that for gases, we can approximate the atomic weight to the density.

OpenStudy (inkyvoyd):

Everything else, I completely forget.

OpenStudy (anonymous):

molar mass = 2* vapour density

OpenStudy (anonymous):

options are 10 20 30 40

OpenStudy (anonymous):

the answer is 30

OpenStudy (inkyvoyd):

what is ab2, and ab4?

OpenStudy (anonymous):

a is an element and b is another element ab2 is its formula

OpenStudy (inkyvoyd):

I don't think I can help you, or maybe i just don't know what you are talking about.. Sorry :(

OpenStudy (inkyvoyd):

Wait, I might have it. Give me a sec.

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

ab2 and ab4 molar mass of ab2=64 molar mass of ab4=96

OpenStudy (unklerhaukus):

Sulfur and Oxygen

OpenStudy (inkyvoyd):

x+y=50 , where x is the amount in moles of ab2, and y is the amount in moles of ab4

OpenStudy (anonymous):

yes it is sulphur and oxygen

OpenStudy (inkyvoyd):

Then, 64y=96x

OpenStudy (inkyvoyd):

Simplify that by dividing by 32 2y=3x

OpenStudy (inkyvoyd):

x+y=50 then, 2x+2y=100, or, 2x+3x=100, so, x=20, and y=30

OpenStudy (inkyvoyd):

No, I got it all wrong. Give me a second.

OpenStudy (inkyvoyd):

(64*c+92*d)/(c+d)=76.8 c+d=50 I thiiink that should be right, although I'm not sure.

OpenStudy (inkyvoyd):

you'll want to solve for d, I think.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i did nt understand that step (64*c+92*d)/(c+d)=76.8

OpenStudy (anonymous):

now i understood plzz go on with ur answer

OpenStudy (inkyvoyd):

Ok. I think it's because I got my equations wrong again.

OpenStudy (inkyvoyd):

Say that c:d= number of ab2 molecules: number of ab4 molecules.

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

Then, 64c+96d=2*38.4 c+d=1, because the concentration of all of the molecules must be 1.

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

Ok, I might make a lot of mistakes. :S

OpenStudy (inkyvoyd):

32d=76.8-64 d=(12.8)/32 c=1-(12.8/32) now, find c*50

OpenStudy (anonymous):

i did nt understand

OpenStudy (inkyvoyd):

Set c to the concentration of ab2 Set d to the concentration of ab4 Then, c+d must be 1, because the solution is completely made up of ab2 and ab4. Now, 64c+96d=2*38.4, solve for c, and d. c and d are concentrations, not actual values, so we must multiply c*50 to get our answer.

OpenStudy (inkyvoyd):

Took me long enough, but I finally have it :D

OpenStudy (anonymous):

ha...haaa thanzzzz

OpenStudy (inkyvoyd):

You got it? :)